zarkam21:

Your dog sleeps on the floor. He has a mass of 14 kg. The coefficient of static friction between him and the floor is 0.3, and the coefficient of kinetic friction is 0.25. a. What is the weight of your dog? b. What is the normal force acting on your dog? c. What is the maximum force of static friction? d. You push horizontally on your dog with a force of 50 N. Does your dog slide? e. Later, you are sliding your dog across the floor at a speed of 2 m/s. What is the force of kinetic friction acting on your dog?

Vocaloid:

a) weight = mg as usual

Vocaloid:

I'm going to be doing some volunteer work later this afternoon but hopefully @sillybilly123 or @Nnesha should come on

zarkam21:

137.2

zarkam21:

oh okay sure no problem

Vocaloid:

b) normal force is equal to the weight since we are on level ground c) max static friction is mu_s * N as usual, make sure you are using the coefficient of static friction d) if 50 > static friction force from c then yes, otherwise no e) Fnet = F_applied - F_friction = ma i'm actually not sure where to go from here but that's the setup at least :S

zarkam21:

would a be a certain unit?

zarkam21:

@Vocaloid

zarkam21:

@Vocaloid

Vocaloid:

units of weight are N

zarkam21:

b)14kg

Vocaloid:

b) normal force is equal to weight not mass

zarkam21:

oh shoot read wrong. so just 137.2N for b

zarkam21:

oh shoot read wrong. so just 137.2N for b

Vocaloid:

good

zarkam21:

ookay so for c it would be

Vocaloid:

mu_s * N

zarkam21:

okay so it would be (14)*

Vocaloid:

mu_s = coefficient of static friction

zarkam21:

0.3=

Vocaloid:

good, so mu_s * N = ?

zarkam21:

is n the kinetic force

Vocaloid:

N is the normal force

zarkam21:

0.3*137.2=41.16

Vocaloid:

good for d) since the applied force > static friction force then yes the dog slides for e) I'm still not 100% on this but I would guess it's just mu_k * N

zarkam21:

N right

Vocaloid:

for units yes

zarkam21:

for e we can just do it whatever you think, because you have been right so far :))))

Vocaloid:

yeah I'd just go with 0.25 * normal force b/c Idk what else one could do here

zarkam21:

0.25 * 137.2=34.3N

Vocaloid:

good